OpenStudy (anonymous):

differentiate and simplify y=x^x-1

OpenStudy (anonymous):

take log of both sides? lny=lnx^x-1

OpenStudy (anonymous):

we have to use log. diff. to find

OpenStudy (anonymous):

x^x*log(x)+x^x ??

zepdrix (zepdrix):

\[\Large \ln y=\ln(x^x-1)\]Hmm that 1 is causing a problem. We should probably move it to the other side before applying the log.\[\Large \ln(y+1)=\ln(x^x)\]

OpenStudy (anonymous):

the answer is x^(x-1)[(1/x)-1-lnx]. just can't get there. the problem is y=x ^(x-1)

zepdrix (zepdrix):

Oh I see :)

zepdrix (zepdrix):

\[\Large \ln y=\ln(x^{x-1})\] Applying our rule of logs:\[\Large \color{teal}{\log(a^b)\quad=\quad b\cdot \log(a)}\]

zepdrix (zepdrix):

gives us:\[\Large \ln y \quad=\quad (x-1)\ln x\]right?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

From this point we can take our derivative, yay!

zepdrix (zepdrix):

\[\Large \color{royalblue}{\left(\ln y\right)'}\quad=\quad \color{royalblue}{(x-1)'}\ln x+(x-1)\color{royalblue}{\left(\ln x\right)'}\]

zepdrix (zepdrix):

So we need to apply the product rule on the right side, yes? Do you understand how to take the derivative of the left side?

OpenStudy (anonymous):

no

zepdrix (zepdrix):

Remember the derivative of ln x?

OpenStudy (anonymous):

dy/dx

zepdrix (zepdrix):

no no no silly :) \[\Large (\ln x)' \quad=\quad\frac{1}{x}\]Remember? :D

OpenStudy (anonymous):

oh, yeah.

zepdrix (zepdrix):

The same thing happens with (ln y)' except we have to remember one important thing. Since we're taking the derivative with `respect to x` of a `non-x variable`, we need to toss a y' onto it. (which is the same as dy/dx).

zepdrix (zepdrix):

\[\Large (\ln y)' \quad=\quad \frac{1}{y}y'\]

OpenStudy (anonymous):

i still do not see the answer in there. can you put the steps together? lny=lnx^(x-1) then...(with the derivatives put in). please

zepdrix (zepdrix):

Ya we still have a ways to go before we get to the answer :) Ok so taking the derivative of both sides gives us this setup:\[\Large \color{royalblue}{\left(\ln y\right)'}\quad=\quad \color{royalblue}{(x-1)'}\ln x+(x-1)\color{royalblue}{\left(\ln x\right)'}\]Calculating the derivative of the left side gave us:\[\Large \color{orangered}{\frac{1}{y}\;y'}\quad=\quad \color{royalblue}{(x-1)'}\ln x+(x-1)\color{royalblue}{\left(\ln x\right)'}\]

zepdrix (zepdrix):

Calculate the derivative of the first blue part on the right, what do you get? :) Derivative of (x-1) = ?

OpenStudy (anonymous):

(1/y)(dy/dx)=...

zepdrix (zepdrix):

derivative of (x-1) = ? +_+

zepdrix (zepdrix):

comeon gal! This is a nice easy one! :) lol

OpenStudy (anonymous):

1

zepdrix (zepdrix):

yay! good good.\[\Large \color{orangered}{\frac{1}{y}\;y'}\quad=\quad \color{orangered}{(1)}\ln x+(x-1)\color{royalblue}{\left(\ln x\right)'}\]

zepdrix (zepdrix):

We discussed the other derivative a lil bit back. So here it is again, just in case you forgot:\[\Large \color{orangered}{\frac{1}{y}\;y'}\quad=\quad \color{orangered}{(1)}\ln x+(x-1)\color{orangered}{\left(\frac{1}{x}\right)}\]

OpenStudy (anonymous):

yep

zepdrix (zepdrix):

We're trying to find y'. So let's multiply both sides by y. The y and 1/y cancel on the left side, giving us:\[\Large y'\quad=\quad y\left(\ln x+\frac{x-1}{x}\right)\]

zepdrix (zepdrix):

We've got just a little bit more work ahead of us.

zepdrix (zepdrix):

We want the `right side` to be written in terms of x. So we need to change the y to x's using the equation we started with.\[\Large \color{teal}{y=x^{x-1}}\]So our derivative function:\[\Large y'\quad=\quad \color{teal}{y}\left(\ln x+\frac{x-1}{x}\right)\]Becomes:\[\Large y'\quad=\quad \color{teal}{x^{x-1}}\left(\ln x+\frac{x-1}{x}\right)\]

OpenStudy (anonymous):

okay, so why is (-) in the answer: x^(x-1)[(1/x)-1-lnx]?

zepdrix (zepdrix):

Did I miss a negative somewhere? Hmm lemme check.

zepdrix (zepdrix):

Hmm the answer shouldn't look like that. It should be:\[\Large y'\quad=\quad x^{x-1}\left(-\frac{1}{x}+1+\ln x\right)\]

zepdrix (zepdrix):

With all the signs reversed inside the brackets ^ like that. Unless your final answer is showing a (-) outside of the brackets as well.

OpenStudy (anonymous):

okay. i will ask teacher tomorrow. thanks! :)

zepdrix (zepdrix):

cool c: